# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse Multi-Baseline Data

### v3

@@ 2.4 Differentiable volume rendering
Here, $\sigma(\mathbf{r}(t))$ denotes the volume density along the camera ray at point $\mathbf{r}(t)$ in the viewing direction $\mathbf{d}$, and $dt$ represents the step size of the ray at each integration step. $T(t)$ is the cumulative transmittance, which indicates the probability that the ray propagates from $t_{n}$ to $t$ without being intercepted. The $T(t)$ is given by the following equation:
@@
Figure 4: Rendering diagram.
To compute the continuous integral, we discretize the ray by sampling it at equidistant depths and apply the integration rules discussed in the literature [31] to estimate the pixel value $C(r)$. This is given by the following formula:
$\hat{C}(\mathbf{r})=\sum\limits_{i=1}^{N}T_{i}(1-\text{exp}(-\sigma_{i}\delta_{i})).$ (3)
@@
$I=\left[\begin{array}[]{ccc}\hat{C}(\mathbf{r}_{0,0})&\dots&\hat{C}(\mathbf{r}_{w-1,0})\\ \vdots&\ddots&\vdots\\ \hat{C}(\mathbf{r}_{0,h-1})&\dots&\hat{C}(\mathbf{r}_{w-1,h-1})\end{array}\right].$ (5)
@@
2.5 Loss design
We use images rendered from the target’s 3D model at different viewpoints to serve as ground truth to compute the reconstruction error. In this paper, we employ the Huber loss function to quantify the error. The Huber loss between the $i$-th viewpoint’s reconstructed image, $I_{r}^{i}$, and the corresponding ground truth image, $I_{g}^{i}$, is defined as:
$\mathcal{L}_{{\rm{huber}}}^{i}({I_{g}^{i}},{I_{r}^{i}},\gamma)=\left\{\begin{array}[]{l}{\textstyle{1\over 2}}{({I_{g}^{i}}-{I_{r}^{i}})^{2}},{\rm{\quad\quad\quad\thinspace\thinspace}}\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|\leq\gamma\\ \gamma\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|-{\textstyle{1\over 2}}{\gamma^{2}}.{\rm{\quad}}\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|>\gamma\end{array}\right.$ (6)
@@ 2.5 Loss design
$\mathcal{L}=\frac{1}{V}\sum\limits_{i=0}^{V-1}{\mathcal{L}_{{\rm{huber}}}^{i}({I_{g}^{i}},{I_{r}^{i}},\gamma)}.$ (7)
3 Experiments
3.1 Experimental Settings
To train CMAR-Net, we construct a civilian vehicle dataset, including simulated sparse multi-baseline SAR 3D pre-images and multi-view optical images. For validation, we perform extensive comparative experiments with baseline methods on the simulated data, including feature space interpolation. We then validate the generalization performance of CMAR-Net using real-world data and conduct ablation studies on the PRP unit and data augmentation strategy. This section provides a detailed overview of the experiment setup.
3.1.1 Dataset and Augmentation
The sparse multi-baseline SAR 3D pre-images in the dataset were generated using the Civilian Vehicle Radar Data Domes (CVDomes), published by the United States Air Force Laboratory[32]. This data includes full-polarization, far-field X-band electromagnetic scattering data for ten civilian vehicles, covering a 360° azimuth and a pitch angle range from 30° to 60°. The viewpoint distribution and parameter settings of the simulation scene are illustrated in Figure 5 and Table 1. For the training set, data from five vehicles—two sedans, two SUVs, and one pickup truck—are selected. Imaging is performed at eight pitch angles between 44.25° and 46°, with a spacing of 0.1875° between each angle. The 360° azimuth data are divided into 72 sub-apertures, each 5° wide and spaced 0.0625° apart. Sub-aperture imaging is then performed, and finally, 8 randomly selected sub-apertures are non-coherently summed to generate the 3D pre-images for training.
Table 1: Simulated Parameters of the Civilian Vehicle Radar Data Domes[32].
@@ 3.1.1 Dataset and Augmentation
Extrapolation extent $\leqslant 0.25^{\circ}$
@@
Additionally, we design a data augmentation strategy to expand our dataset. As shown in Figure 6, the augmentation involves three operations: scaling $S$, rotation $R$, and translation $T$. For each pair of input raw 3D pre-images, a geometric transformation is applied, which is also performed on the corresponding 3D digital model. This transformation is then reflected in the rendered images. This strategy increases the sampling density of the data space, which enhances the network’s generalization.
Figure 5: Dataset construction.
Figure 6: Data augmentation strategy.
3.1.2 Implementation
In the pre-imaging stage, the imaging scene is set with dimensions of $3.2m\times 3.2m\times 6.4m$ and a spatial resolution of $0.05m$, resulting in a data size of $64\times 64\times 128$. The pre-imaging data is then normalized to a dynamic range of [0, 1] before being input into the network. For the differentiable volume rendering module, the rendered image size is set to 256 $\times$ 256, with the camera positioned 7m from the scene center. The camera is fixed to point toward the center from 8 predetermined angles.
@@ 3.1.4 Evaluation Metrics
We convert the vehicle’s 3D digital model into volume data with the same spatial extent and resolution as the imaging settings, using it as the ground truth. The imaging quality of our method is then evaluated using full-reference metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Intersection over Union (IoU), and Cross-Entropy (CE).
3.2 Results for Simulated Data
3.2.1 Pre-imaging and Reconstruction Results
First, we provide an overview of the 3D reconstruction performance of CMAR-Net. Figure 7 illustrates the pre-imaging results fed into the network and the reconstruction results produced by the CMAR-Net. As shown in the figure, CMAR-Net significantly enhances the reconstruction of the vehicle’s surface, resulting in a reconstructed image with visually distinguishable vehicle features that closely resemble the ground truth (GT).
Figure 7: Overview of 3D reconstruction performance of CMAR-Net. Pre-imaging results (first row), CMAR-Net reconstructions (second row) and ground truth (third row). Imaging at an SNR of 30dB and 8 sub-apertures.
3.2.2 Comparison with Baseline Methods
Next, we compare CMAR-Net with baseline methods. Figure 8 presents the imaging results of different methods at an SNR of 30dB and 8 sub-apertures. Compared to the pre-imaging results in Figure 7, the BP method preserves the strong scattering points of the vehicle body by non-coherently integrating the maximum amplitude values from sub-aperture images. For instance, the bucket structure of Pickup are prominently improved. However, the imaging quality is degraded by high side lobes, as evidenced by the distorted front and windshield of the MPV and Camry. The CS method mitigates the impact of side lobes by optimizing sub-aperture imaging. However, the imaging results reveal significant structural omissions, attributed to the sparsity assumptions and the configuration of the optimization parameters.
Figure 8: Imaging results of different methods. Each column corresponds to an imaging method, and each row represents a vehicle type. Imaging at an SNR of 30dB and 8 sub-apertures.
To evaluate the reconstruction performance of CMAR-Net under different conditions, we conduct experiments where the SNR decreased from 25 dB to 5 dB, and the number of sub-apertures reduced from 12 to 4. Figures 9 to 13 present the imaging results for all experimental conditions.
As the SNR decreases and the number of sub-apertures reduces, the imaging results of the BP and CS methods show more side lobes, leading to reduced vehicle recognizability. SACNet and UNet3D demonstrate strong robustness under deteriorating imaging conditions, providing consistent reconstruction results. After integrating the cross-modal supervision, their performance remains stable. In contrast, CMAR-Net consistently delivers the best imaging performance across all vehicle types. Even under the most challenging conditions (5 dB SNR and 4 sub-apertures), CMAR-Net achieves excellent reconstruction, while the performance of other methods significantly deteriorates.
Figure 9: Imaging results of different methods at an SNR of 25dB and 12 sub-apertures.
Figure 10: Imaging results of different methods at an SNR of 20dB and 10 sub-apertures.
Figure 11: Imaging results of different methods at an SNR of 15dB and 8 sub-apertures.
Figure 12: Imaging results of different methods at an SNR of 10dB and 6 sub-apertures.
Figure 13: Imaging results of different methods at an SNR of 5dB and 4 sub-apertures.
3.2.3 Quantitative comparison
We calculate the quality metrics under different data conditions, with the results presented in Tables 2 to 5. Figures 14 visualizes these results using curve plots. The data reveal that all existing methods perform worse across all metrics compared to DL-based methods with cross-modal supervision, with particularly notable differences in the IoU and CE metrics. This disparity is largely due to modality differences between the output images and the ground truth data. And among the proposed DL-based methods, CMAR-Net consistently achieves the best performance across all metrics. The only exception occurs under the condition of 4 sub-apertures and an SNR of 5 dB, where CMAR-Net’s SSIM is slightly lower than the method using UNet3D as the backbone. We believe this is due to the UNet3D’s tendency to output higher pixel values (closer to 1). However, this behavior in fact leads to a loss of structural information, exhibiting poor generalization of the network, a conclusion that will be confirmed in real-world experiments.
Table 2: IoU $\uparrow$ comparison across all methods on simulated data.
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP[35] 0.051 0.055 0.045 0.077 0.067 0.051 0.064 0.047 0.055
CS[34] 0.032 0.036 0.031 0.059 0.025 0.031 0.046 0.027 0.037
SACNet[23] 0.078 0.081 0.069 0.079 0.075 0.073 0.075 0.078 0.074
UNet3D[22] 0.031 0.030 0.041 0.027 0.041 0.030 0.038 0.034 0.046
SACNet+CM 0.653 0.744 0.269 0.717 0.755 0.434 0.558 0.462 0.592
UNet3D+CM 0.639 0.638 0.601 0.639 0.604 0.553 0.593 0.586 0.545
CMAR-Net(Ours) 0.750 0.744 0.713 0.745 0.726 0.689 0.694 0.670 0.626
Table 3: CE $\downarrow$ comparison across all methods on simulated data.
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP 1.624 1.685 1.637 1.625 1.640 1.636 1.656 1.629 1.654
CS 1.634 1.647 1.637 1.622 1.631 1.632 1.643 1.640 1.638
SACNet 1.307 1.294 1.091 1.357 1.196 1.263 1.283 1.332 0.992
UNet3D 0.491 0.503 0.593 0.511 0.545 0.565 0.526 0.533 0.607
SACNet+CM 0.205 0.171 0.530 0.215 0.153 0.371 0.464 0.498 0.387
UNet3D+CM 0.167 0.174 0.187 0.154 0.195 0.171 0.154 0.179 0.186
CMAR-Net(Ours) 0.099 0.108 0.096 0.112 0.109 0.104 0.119 0.111 0.139
@@ Methods
BP 8.555 8.393 8.523 8.557 8.518 8.532 8.473 8.553 8.475
CS 8.538 8.502 8.530 8.573 8.548 8.549 8.511 8.525 8.527
SACNet 8.798 8.796 8.799 8.734 8.801 8.723 8.774 8.744 8.815
UNet3D 9.220 9.193 9.044 9.164 9.150 9.028 9.176 9.150 9.034
SACNet+CM 13.213 14.540 10.477 13.739 14.711 11.672 10.470 10.648 11.417
UNet3D+CM 13.183 13.143 12.964 13.409 12.806 13.157 13.333 12.971 12.897
CMAR-Net(Ours) 16.446 16.212 16.236 16.107 15.822 15.774 15.384 15.448 14.329
@@
BP 0.612 0.574 0.612 0.580 0.584 0.604 0.574 0.605 0.584
CS 0.626 0.603 0.628 0.591 0.637 0.628 0.589 0.626 0.606
SACNet 0.626 0.613 0.595 0.615 0.608 0.588 0.585 0.574 0.486
UNet3D 0.447 0.452 0.484 0.460 0.477 0.476 0.480 0.457 0.503
SACNet+CM 0.667 0.718 0.695 0.696 0.724 0.698 0.486 0.536 0.559
UNet3D+CM 0.763 0.762 0.750 0.774 0.744 0.757 0.766 0.753 0.746
CMAR-Net(Ours) 0.820 0.811 0.796 0.805 0.796 0.776 0.782 0.778 0.737
@@
3.2.4 Feature space interpolation
Figure 15 shows the results of the feature space interpolation experiment. In this experiment, we perform linear interpolation between feature vectors extracted by the encoder structure in CMAR-Net, and observe how these interpolated features affect the generated output. The green boxes in the figure represent the output vehicle images of the training type. The images between each pair of green boxes represent the output results of the interpolated feature space. As shown in the figure, linear interpolation between two feature vectors leads to smooth and meaningful nonlinear transitions in the image space[36, 37]. Notably, the variations in body height and tail features of the interpolated images do not align with any specific vehicle type from the training data. This indicates that the network preserves distinct vehicle features during interpolation, rather than averaging them into an ”average” representation[38]. We believe this is a superior property of the reconstruction network, as it highlights the model’s ability to generate diverse outputs. This shows the model’s flexibility and capability to produce realistic and varied vehicle reconstructions.

### v4

@@ II-E Differentiable volume rendering
Here, $\sigma(\mathbf{r}(t))$ and ${\bf{c}}({\bf{r}}(t),{\bf{d}})$ represent the volume density and color at point $\mathbf{r}(t)$ along the camera ray with viewing direction $\mathbf{d}$, and $dt$ represents the step size of the ray at each integration step. $T(t)$ is the cumulative transmittance, which indicates the probability that the ray propagates from $t_{n}$ to $t_{f}$ without being intercepted. The $T(t)$ is given by the following equation:
@@
To compute the continuous integral, we discretize the ray by sampling it at equidistant depths and apply the integration rules discussed in the literature [32] to estimate the pixel value $C(\mathbf{r})$. This is given by the following formula:
$\hat{C}(\mathbf{r})=\sum\limits_{i=1}^{N}T_{i}(1-\text{exp}(-\sigma_{i}\delta_{i})){{\bf{c}}_{i}},$ (3)
@@
$I=\left[\begin{array}[]{ccc}\hat{C}(\mathbf{r}_{0,0})&\dots&\hat{C}(\mathbf{r}_{w-1,0})\\ \vdots&\ddots&\vdots\\ \hat{C}(\mathbf{r}_{0,h-1})&\dots&\hat{C}(\mathbf{r}_{w-1,h-1})\end{array}\right],$ (5)
@@
In volume rendering models, each voxel location within 3D space typically encodes two fundamental attributes: volume density $\sigma$ and color $\bf{c}$. The density parameter quantifies a spatial position’s light attenuation characteristics, where higher density values indicate greater probability of light occlusion. In several existing 3D reconstruction methods [33, 34, 35], it is commonly assumed that regions with higher density are more likely to correspond to actual physical structures. As a result, density is often used as a proxy for the probability of object presence in space, and is subsequently employed to generate the 3D structural image of the object.
The electromagnetic nature of SAR makes it challenging to model color information. In our work, we focus on reconstructing the object’s geometric structure by using the volumetric density parameter as a proxy for the spatial occupancy probability. We simplify the rendering process by treating the color term $\bf{c}$ as a constant, effectively ignoring the influence of color. Meanwhile, to ensure consistent supervision, the optical images of vehicles used for training are preprocessed into binary masks that retain only geometric shape information. In these masks, pixels with a value of 1 indicate the presence of vehicle structures and serve as strong supervision signals for object existence in 3D space. This design encourages the network to assign higher density values to corresponding regions. Through sufficient training, the network learns the structural distribution of the target from these supervisory images and ultimately produces 3D reconstructions of vehicles that are spatially well-aligned with the optical supervision.
II-F Loss design
We use mask images rendered from 3D vehicle models viewed from different angles as ground truth to compute the reconstruction error. During the training process, we employ the Huber loss function to quantify the error. The Huber loss between the $i$-th viewpoint’s reconstructed image, $I_{\rm{r}}^{i}$, and the corresponding ground truth image, $I_{\rm{g}}^{i}$, is defined as:
$\mathcal{L}_{{\rm{huber}}}^{i}({I_{\rm{g}}^{i}},{I_{\rm{r}}^{i}},\gamma)=\begin{cases}\frac{1}{2}\left({I_{\rm{g}}^{i}}-{I_{\rm{r}}^{i}}\right)^{2}&\text{if }\left|{I_{\rm{g}}^{i}}-{I_{\rm{r}}^{i}}\right|\leq\gamma,\\ \gamma\left|{I_{\rm{g}}^{i}}-{I_{\rm{r}}^{i}}\right|-\frac{1}{2}\gamma^{2}&\text{if }\left|{I_{\rm{g}}^{i}}-{I_{\rm{r}}^{i}}\right|>\gamma.\end{cases}$ (6)
@@ II-F Loss design
$\mathcal{L}=\frac{1}{V}\sum\limits_{i=0}^{V-1}{\mathcal{L}_{{\rm{huber}}}^{i}({I_{\rm{g}}^{i}},{I_{\rm{r}}^{i}},\gamma)}.$ (7)
Figure 6: Dataset construction.
Figure 7: Data augmentation strategy.
III Experiments
III-A Experimental Settings
To train CMAR-Net, we construct a civilian vehicle dataset, including simulated sparse-aspect multi-baseline SAR 3D images and multi-view optical images. For validation, we perform extensive comparative experiments with baseline methods on the simulated data, including feature space interpolation. We then validate the generalization performance of CMAR-Net using real-world data and conduct ablation studies on the PRP unit and data augmentation strategy. This section provides a detailed overview of the experiment setup.
III-A1 Dataset and Augmentation
The sparse-aspect multi-baseline SAR 3D images in the dataset were generated using the Civilian Vehicle Radar Data Domes (CVDomes), published by the United States Air Force Laboratory[29]. This data includes full-polarization, far-field X-band electromagnetic scattering data for ten civilian vehicles, covering a 360° azimuth and a pitch angle range from 30° to 60°. The viewpoint distribution and parameter settings of the simulation scene are illustrated in Figure 6 and Table I. For the training set, data from five vehicles—two sedans, two SUVs, and one pickup truck—are selected. Imaging is performed at eight pitch angles between 44.25° and 46°, with a spacing of 0.1875° between each angle. The 360° azimuth data are divided into 72 sub-aspects, each 5° wide and spaced 0.0625° apart. Sub-aspect imaging is then performed, and finally, 8 randomly selected sub-aspects are non-coherently summed to generate the 3D SAR images for training.
TABLE I: Simulated Parameters of the Civilian Vehicle Radar Data Domes and Training data setting[29].
@@ III-A1 Dataset and Augmentation
Extrapolation extent $\leq 0.25^{\circ}$
@@
Sub-aspect angle extent $5^{\circ}$
Sub-aspect number $72$
Elevation range for training $44.25^{\circ}-46^{\circ}$
@@
Additionally, we design a data augmentation strategy to expand our dataset. As shown in Figure 7, the augmentation involves three operations: scaling $S$, rotation $R$, and translation $T$. For each pair of input raw 3D SAR image, a geometric transformation is applied, which is also performed on the corresponding 3D digital model. This transformation is then reflected in the rendered images. This strategy increases the sampling density of the data space, which enhances the network’s generalization.
Figure 8: Overview of 3D reconstruction performance of CMAR-Net. BP results (1st row), CMAR-Net reconstructions (2nd-4th rows) and ground truth (5th row). Imaging at an SNR of 30dB and 8 sub-aspects.
Figure 9: Imaging results of different methods. Each column corresponds to an imaging method, and each row represents a vehicle type. Imaging at an SNR of 30dB and 8 sub-aspects.
III-A2 Implementation
In the multi-aspect 3D SAR imaging stage, the imaging scene is set with dimensions of $3.2m\times 3.2m\times 6.4m$ and a spatial resolution of $0.05m$, resulting in a data size of $64\times 64\times 128$. The 3D SAR data is then normalized to a dynamic range of [0, 1] before being input into the network. For the differentiable volume rendering module, the rendered image size is set to 256 $\times$ 256, with the camera positioned 7m from the scene center. The camera is fixed to point toward the center from 8 predetermined angles.
@@ III-A4 Evaluation Metrics
We convert the vehicle’s 3D digital model into volume data with the same spatial extent and resolution as the imaging settings, using it as the ground truth. The imaging quality of our method is then evaluated using full-reference metrics, including Peak Signal-to-Noise Ratio (PSNR)[38], Structural Similarity Index Measure (SSIM)[38], Intersection over Union (IoU)[3], and Cross-Entropy (CE)[3].
III-B Results for Simulated Data
III-B1 3D Reconstruction Results of CMAR-Net
Figure 8 presents the 3D reconstruction results of CMAR-Net under an SNR of 30 dB and with 8 sub-aspects, alongside the corresponding BP imaging results and 3D model visualizations. To comprehensively illustrate the reconstruction quality, the CMAR-Net output is visualized from three viewing angles with (azimuth, elevation) parameters of (0°, 17°), (30°, 17°), and (135°, 17°). For comparison, the BP imaging results and its corresponding 3D model are shown using the (0°, 17°) viewpoint. Unless otherwise specified, this viewpoint is used by default in all subsequent experimental visualizations. As shown in the figure, the BP method, limited by the use of only 8 sub-aspect observations, produces significant sidelobe artifacts and incomplete vehicle structures. In contrast, CMAR-Net substantially enhances the reconstruction of the vehicle surface, successfully recovering visually distinguishable geometric features. The resulting shape closely matches the corresponding 3D model.
Figure 10: Imaging results of different methods at an SNR of 25dB and 12 sub-aspects.
Figure 11: Imaging results of different methods at an SNR of 20dB and 10 sub-aspects.
Figure 12: Imaging results of different methods at an SNR of 15dB and 8 sub-aspects.
Figure 13: Imaging results of different methods at an SNR of 10dB and 6 sub-aspects.
Figure 14: Imaging results of different methods at an SNR of 5dB and 4 sub-aspects.
III-B2 Comparison with Baseline Methods
@@ Methods
BP[39] 0.051/1.624 0.055/1.685 0.045/1.637 0.077/1.625 0.067/1.640 0.051/1.636 0.064/1.656 0.047/1.629 0.055/1.654
CS[37] 0.032/1.634 0.036/1.647 0.031/1.637 0.059/1.622 0.025/1.631 0.031/1.632 0.046/ 1.643 0.027/1.640 0.037/1.638
SACNet[3] 0.078/1.307 0.081/1.294 0.069/1.091 0.079/1.357 0.075/1.196 0.073/1.263 0.075/1.283 0.078/1.332 0.074/0.992
UNet3D[23] 0.031/0.491 0.030/0.503 0.041/0.593 0.027/0.511 0.041/0.545 0.030/ 0.565 0.038/0.526 0.034/0.533 0.046/ 0.607
SACNet+CM 0.653/0.205 0.744/0.171 0.269/0.530 0.717/0.215 0.755/0.153 0.434/0.371 0.558/0.464 0.462/0.498 0.592/0.387
UNet3D+CM 0.639/0.167 0.638/0.174 0.601/0.187 0.639/0.154 0.604/0.195 0.553/0.171 0.593/0.154 0.586/0.179 0.545/ 0.186
CMAR-Net(Ours) 0.750/0.099 0.744/0.108 0.713/0.096 0.745/0.112 0.726/0.109 0.689/0.104 0.694/0.119 0.670/0.111 0.626/0.139
@@
BP 0.612/8.56 0.574/8.39 0.612/8.52 0.580/8.56 0.584/8.51 0.604/8.53 0.574/8.47 0.605/8.55 0.584/8.48
CS 0.626/8.54 0.603/8.50 0.628/8.53 0.591/8.57 0.637/8.55 0.628/8.55 0.589/8.51 0.626/8.53 0.606/8.53
SACNet 0.626/8.80 0.613/8.80 0.595/8.80 0.615/8.73 0.608/8.80 0.588/8.72 0.585/8.77 0.574/8.74 0.486/8.82
UNet3D 0.447/9.22 0.452/9.04 0.484/9.04 0.460/9.16 0.477/9.15 0.476/9.03 0.480/9.18 0.457/9.15 0.503/9.03
SACNet+CM 0.667/13.21 0.718/14.54 0.695/10.48 0.696/13.74 0.724/14.71 0.698/11.67 0.486/10.47 0.536/10.65 0.559/11.42
UNet3D+CM 0.763/13.18 0.762/13.14 0.750/12.96 0.774/13.41 0.744/12.81 0.757/13.16 0.766/13.33 0.753/12.97 0.746/12.90
CMAR-Net(Ours) 0.820/16.45 0.811/16.21 0.796/16.24 0.805/16.11 0.796/15.82 0.776/15.77 0.782/15.38 0.778/15.45 0.737/14.33
@@
We compare CMAR-Net with baseline methods. Figure 9 presents the imaging results of different methods at an SNR of 30dB and 8 sub-aspects. The BP method preserves the strong scattering points of the vehicle body by non-coherently integrating the maximum amplitude values from sub-aspect images. For instance, the bucket structure of Pickup are prominently improved. However, the imaging quality is degraded by high side lobes, as evidenced by the distorted front and windshield of the MPV and Camry. The CS method mitigates the impact of side lobes by optimizing sub-aspect imaging. However, the imaging results reveal significant structural omissions, attributed to the sparsity assumptions and the configuration of the optimization parameters.
To evaluate the reconstruction performance of CMAR-Net under different conditions, we conduct experiments where the SNR decreased from 25 dB to 5 dB, and the number of sub-aspects reduced from 12 to 4. Figures 10 to 14 present the imaging results for all experimental conditions.
As the SNR decreases and the number of sub-aspects reduces, the imaging results of the BP and CS methods show more side lobes, leading to reduced vehicle recognizability. SACNet and UNet3D demonstrate strong robustness under deteriorating imaging conditions, providing consistent reconstruction results. After integrating the cross-modal supervision, their performance remains stable. In contrast, CMAR-Net consistently delivers the best imaging performance across all vehicle types. Even under the most challenging conditions (5 dB SNR and 4 sub-aspects), CMAR-Net achieves excellent reconstruction, while the performance of other methods significantly deteriorates.
III-B3 Quantitative comparison
We calculate the quality metrics under different data conditions, with the results presented in Tables II to III. Figures 15 visualizes these results using curve plots. The data reveal that all existing methods perform worse across all metrics compared to DL-based methods with cross-modal supervision, with particularly notable differences in the IoU and CE metrics. This disparity is largely due to modality differences between the output images and the ground truth data. And among the proposed DL-based methods, CMAR-Net consistently achieves the best performance across all metrics. The only exception occurs under the condition of 4 sub-aspects and an SNR of 5 dB, where CMAR-Net’s SSIM is slightly lower than the method using UNet3D as the backbone. We believe this is due to the UNet3D’s tendency to output higher pixel values (closer to 1). However, this behavior in fact leads to a loss of structural information, exhibiting poor generalization of the network, a conclusion that will be confirmed in real-world experiments.
III-B4 Feature space interpolation
Figure 16 shows the results of the feature space interpolation experiment. In this experiment, we perform linear interpolation between feature vectors extracted by the encoder structure in CMAR-Net, and observe how these interpolated features affect the generated output. The green boxes in the figure represent the output vehicle images of the training type. The images between each pair of green boxes represent the output results of the interpolated feature space. As shown in the figure, linear interpolation between two feature vectors leads to smooth and meaningful nonlinear transitions in the image space[40, 41, 42]. Notably, the variations in body height and tail features of the interpolated images do not align with any specific vehicle type from the training data. This indicates that the network preserves distinct vehicle features during interpolation, rather than averaging them into an ”average” representation[30]. We believe this is a superior property of the reconstruction network, as it highlights the model’s ability to generate diverse outputs. This shows the model’s flexibility and capability to produce realistic and varied vehicle reconstructions.